OpenStudy (anonymous):

Solve log6x^8-log6x^5=31

OpenStudy (solomonzelman):

at first use \(\Large\color{black}{ \bf log(a^{b})=b~log(a) }\)

OpenStudy (solomonzelman):

re-write using the above rule \(\Large\color{black}{1)~~~~ \bf log~(6x~^{8})~=~? }\) \(\Large\color{black}{2)~~~~ \bf log~(6x~^{5})~=~? }\) the re-write the problem again.

OpenStudy (solomonzelman):

then re-write...

OpenStudy (tkhunny):

Is it this? \(log_{6}\left(x^{8}\right)\)

OpenStudy (anonymous):

yes , well the x^8 wasn't in "( )" but yes

OpenStudy (solomonzelman):

\(\Large\color{black}{\bf log_6x^8-log_6x^5=31 }\)

OpenStudy (anonymous):

@SolomonZelman wouldn't it rewrite to \[8 \log_{6} (x)-5\log_{g}(x) =31\]

OpenStudy (solomonzelman):

YES! \(\Large\color{black}{\bf 8~log_6x-5~log_6x=31 }\)

OpenStudy (solomonzelman):

Which simplifies on the left side to, \(\Large\color{black}{\bf 3~log_6x=31 }\)

OpenStudy (solomonzelman):

can you finish the problem or need more help?

OpenStudy (anonymous):

then you divide both sides by 3 correct?

OpenStudy (solomonzelman):

I am lazy and have no calc.... \(\Large\color{black}{\bf 3log_6x=31 \times log_66 }\) \(\Large\color{black}{\bf log_6x^{3}= log_66^{31} }\) \(\Large\color{black}{\bf x^{3}= 6^{31} }\)

OpenStudy (anonymous):

to get x=\[6 ^{10\frac{ 1 }{ 3 }}\]

OpenStudy (solomonzelman):

See what I did in my last post ?

OpenStudy (anonymous):

yeah, but i don't seem to understand how you got that.If you don't mind could explain that please?

OpenStudy (tkhunny):

\(\log^{b}(a) = c \iff b^{c} = a\)

OpenStudy (solomonzelman):

Yes, I multiplied the left side times magic 1, (which is log(base 6) ( 6 ) )

OpenStudy (solomonzelman):

I multiplied the right side, not the left side, ... sorry.

OpenStudy (anonymous):

ohh i understand now thanks

OpenStudy (solomonzelman):

You welcome !

OpenStudy (anonymous):

If you don't mind could you help me with one more problem please?

OpenStudy (tkhunny):

\(\log_{b}(a) = c\iff b^{c} = a\) Not sure how that b managed to get up in the attic.